OpenStudy (daenio):

If the measurements of the radius of a piston are normally distributed with a mean of 1.85 inches and a standard deviation of 0.25 inches, determine the probability that the radius of a piston randomly chosen from a production assembly line has a radius between 1.7 and 2.1 inches. Question 4 options: a) 2.67% b) 56.7% c) 15.83% d) 11.56%

OpenStudy (daenio):

@lalaly @blockcolder @ash2326 @ZhangYan

OpenStudy (blockcolder):

The z-score of 1.7 is (1.7-1.85)/0.25=-0.6 while the z-score of 2.1 is (2.1-1.85)/0.25=1. The probability that the data is to the left of 1.7 is -0.6 is 0.2743 while the probability that the data is to the left of 2.1 is 0.8413. (Consult a table for this.) Thus, the probability that the data is between the two is 0.8413-0.2743=0.567=56.7%.

OpenStudy (daenio):

Thank you!